# Cyclic Quantum Annealing: Searching for Deep Low-Energy States in 5000-Qubit Spin Glass

### v1

@@ 6 Methods
This section outlines the details of the cyclic annealing algorithm. The algorithm is implemented on D-Wave’s Advantage 4.1 quantum processor, equipped with 5627 working qubits and 40279 couplers.
@@
The initial qubit state for the first cycle can be any state. Interestingly, starting with a high-energy state does not hurt the algorithm’s efficiency, as it rapidly converges to a low-energy state after the first cycle. An important aspect of the algorithm involves specifying the duration for each segment of the path. Table 1 presents a typical path with corresponding time allocations, as used in the performance study described in Section 3.
@@
$(b_{z},b_{x})$ time ($\mu s$)
$(0,0)$ 00
$(0.03,0)$ $0.1$
$(0.03,8.04)$ $0.6$
$(0,0)$ $300$
During the first step of the algorithm, all eigenstates are bitstring states, as the non-commuting driving Hamiltonian (Eq. 4) is not yet turned on. The evolution in this part can be quick. The second step, characterized by a finite energy gap above the ground state, can also be executed rapidly without consuming substantial annealing time. However, the third step, which goes into the spin-glass phase and encounters numerous Landau-Zener transitions, necessitates a slower pace.
For each cycle, the initial state is identical to the reference state of that cycle. These terms are used interchangeably, depending on the context. It is noteworthy that the cycle’s shape is not confined to a triangular or trapezoidal shape. An intriguing and open question remains on how different cycle shapes influence the annealing performance and the quest for an optimal one.
We are grateful to Mohammad Amin, Kevin Chern, Pau Farré, Fiona Hanington, Emile Hoskinson, Andrew King, Jack Raymond, Hanteng Wang and Hsiu-Chung Yeh for useful discussions. This work was supported by the NSF grants DMR-2037654 and DMR-2338819.

### v2

@@ VI Methods
This section outlines the details of the cyclic quantum annealing algorithm and simulations. The algorithm is implemented on D-Wave’s Advantage 4.1 quantum processor, which features 5,627 functional qubits and 40,279 couplers. We utilized all available qubits and couplers in our studies.
The problem set in the Simulations section consists of 50 random realizations of spin glass configurations: for Problem 0 to problem 9, $J_{ij}$ is randomly chosen as ±1; for problem 10 to problem 19, $J_{ij}$ is randomly chosen as ±1, ±2; for problem 20 to problem 29, $J_{ij}$ is randomly chosen as ±1, ±2, ±3, ±4; for problem 30 to problem 39, $J_{ij}$ is randomly chosen as ±1, ±2, ±3, ±4 … ±8; for problem 40 to problem 49, $J_{ij}$ is randomly chosen as ±1, ±2, ±3, ±4 … ±16. Given that the underlying topology closely resembles a 3D lattice, these problems represent 3D spin glass systems, which are known to be NP-hard. We normalize the problem to have average $|J_{ij}|$=1 before the simulations.
@@
For the cyclic annealing, the initial qubit state for the first cycle can be arbitrary $z$ polarized product state. Interestingly, starting with a high-energy state does not hurt the algorithm’s efficiency, as it rapidly converges to a low-energy state after the first cycle. An important aspect of the algorithm involves specifying the duration of each segment of the path. Table 1 presents a typical path with corresponding time allocations, as used in the performance study described in the Simulations Section.
@@
time ($\mu s$) $(b_{z},b_{x})$
00 $(0,0)$
$0.1$ $(0.03,0)$
$0.6$ $(0.03,8.04)$
$300$ $(0,0)$
During the first step of the algorithm, all eigenstates are bitstring states, as the non-commuting driving Hamiltonian (Eq. 4) is not yet turned on. The evolution in this part can be fast. The second step, characterized by a finite energy gap above the ground state, can also be executed rapidly without consuming substantial annealing time. However, the third step, which goes into the spin-glass phase and encounters numerous Landau-Zener transitions, necessitates a slower pace. It is recommended to set the annealing time within the range of 100 to 1000 $\mu$s. In our experiments, we used 300 $\mu$s in the Simulations section and 500 $\mu$s in the subsequent sections.
It is noteworthy that the cycle’s shape is not confined to a triangular or trapezoidal shape. An intriguing and open question remains on how different cycle shapes influence the annealing performance and the quest for an optimal one.
VII acknowledgments
We are grateful to Mohammad Amin, Kevin Chern, Pau Farré, Fiona Hanington, Emile Hoskinson, Andrew King, Jack Raymond, Hanteng Wang, and Hsiu-Chung Yeh for useful discussions. This work was supported by the NSF grants DMR-2037654 and DMR-2338819.
VIII Data availability
All data analyzed in this study are available from the corresponding author upon reasonable request.